OpenStudy (karinamtz0213):

The line given by the equation y= -3x+1 goes through the points (1,a) and (b,4). 2a-3b= ? a. -2 b. -1 c. 0 d. 1 e. Cannot be determined from given information

OpenStudy (anonymous):

if you know \(x\) then you know \(y\) and vice versa

OpenStudy (anonymous):

if \(x=1\) then \(y=-3\times 1+1=-2\) so \((1,-2)\) is on the graph, i.e. \(a=-2\)

OpenStudy (anonymous):

It is simple if you graph the equation. Look where x = 1, y = -2. And where y = 4, x = -1 So then you substitute a and b in to the second equation. a = -2 b = -1 2a - 3b = 2(-2) - 3(-1) = -4 - (-3) = -4 + 3 = -1 So -1 is the answer.